OpenStudy (marigirl):

If the mean of the first 10 numbers is 11.9 and the mean of the second 10 numbers is 8.0. find the mean of all 20 numbers.

OpenStudy (anonymous):

9.95

OpenStudy (marigirl):

how did u get that?

OpenStudy (anonymous):

I would do 11.9+8 divided by 2=19.9 divided by 2=9.95

OpenStudy (anonymous):

11.9*10 gives sum of first 10 8*10 gives the sum of second 10 sum of 20 numbers are 119+80=199 mean=199/20=9.95

OpenStudy (anonymous):

@Loujoelou is also right

OpenStudy (anonymous):

but only when size of the first set 10 number and size of second set 10 numbers are equal...it will not hold good when first set has 11 numbers and second set has 8 numbers

OpenStudy (marigirl):

Thank you to both of you.. why cant i give best answer to two people :(

OpenStudy (marigirl):

thank you @Loujoelou

OpenStudy (anonymous):

No prob. :)